# EXCOGITO, an extensible coarse-graining toolbox for the investigation of biomolecules by means of low-resolution representations

### v1

@@ 2.2 Mapping entropy and related EXCOGITO tools
EXCOGITO enables the identification of the maximally informative CG representation of a system via the optimize and optimize_kl tools, which respectively build on the equilibrium and KL mapping entropy estimators reported in Eqs. 13 and  14. Both protocols minimize $S_{map}$ by relying on a Monte Carlo simulated annealing (SA) approach, gradually pushing a stochastic exploration of the mapping space of the system to visit CG representations characterized by a low information loss, see Sec.2.4 for all technical details. More specifically, starting from an initial selection of $N$ sites of the macromolecule, optimize and optimize_kl perform a sequence of Monte Carlo moves that, at each step, propose a swap between a retained and neglected atom in the CG mapping, hence working at a fixed degree of coarse-graining. The moves are accepted according to a Metropolis-like criterion that employs $S_{map}$ as a cost function, and in which the associated “temperature” is exponentially decreased in the course of the simulation, driving the sampling, after an initial transient, to converge towards a local minimum of the mapping entropy. A single SA run performed with one of the two methods thus enables the identification of one of the sought-for maximally informative CG representations that can be employed to describe the system of interest.
The next necessary step in this analysis is to account for the fact that the manifold of solutions to the optimization problem can have a rather complex structure. Indeed, given the intricacy of the network of interactions among the system’s microscopic constituents, it is reasonable to expect a rugged landscape of information loss throughout the mapping space, exhibiting a whole ensemble of more or less degenerate local minima, either living in relatively flat basins or being widely separated by high $S_{map}$ barriers. None of these minima can a priori be preferred over another; rather, in order to gather a full picture of the link between resolution reduction and information content, one needs to simultaneously consider a pool of solutions minimising the mapping entropy. As we will discuss in the following, it is precisely the pattern that emerges from the analysis of the whole ensemble of such optimized solutions that enables the extraction of nontrivial information about the system and its function.
With reference to the software, as in the case of the unbiased exploration of the mapping space performed by random and random_kl, we note that this minimization protocol is independently implemented in two EXCOGITO tools, optimize and optimize_kl, which respectively rely on the equilibrium and KL mapping entropy estimators $\tilde{S}_{map}$ and $\hat{S}_{map}$ in Eqs. 13 and  14.
@@ 2.3 Metric in the coarse-grained mapping space and related EXCOGITO tools
Equations 12 and 14 allow us to calculate the mapping entropy associated to a given mapping, and their optimisation leads to the identification of mappings that entail the largest amount of information about the system for a given number of retained atoms. It is also instructive, however, to broaden the perspective on mappings themselves, and to investigate their properties from a purely structural perspective.
More specifically, given a mapping as the selection of a particular subgroup of elements (the retained atoms) from a set (the whole molecule), we can ask ourselves questions about the total number of mappings, the amount of mappings sharing the same qualitative and quantitative features, and the relationship that exists between mapping groups with given features and the structural properties of the system on which the selection takes place. These questions have been addressed in various recent works 35, 34, 36; here, we only report those results obtained by some of us 34 through the application methods implemented in EXCOGITO.
In order to assess even the simplest properties of the mapping space associated to a given protein, one needs to define basic quantitative instruments, for example to determine how different a given mapping is from another. To this end, some of us introduced a notion of norm, cosine, and distance between coarse-grained mappings, which only make use of the structural features of the biomolecule of interest. More specifically, we define the scalar product $\langle{\bf M},{{\bf M}^{\prime}}\rangle$ between two mappings $\bf M$} and $\bf M$$^{\prime}$ as:
@@
where $\chi_{{\bf M},j}$ is the mapping function as defined in Eq.7, $r_{ij}$ is the distance between atoms $i$ and $j$, and $\sigma$ is a parameter that tunes the amplitude of the Gaussian employed to calculate the coupling $J_{ij}=e^{-r^{2}_{ij}/4\sigma^{2}}$ between them. In the work in which it was introduced, this parameter was set to $\sigma=0.18$ nm, however its value can be tuned to the specific application.
From Eq. 17 we can calculate the norm of a mapping as
@@
In Ref. 34 the norm and the distance are rescaled by a function of the atomistic coordination number:
@@
calculated over a specific molecular configuration (e.g. the initial frame of an MD trajectory, or the frame closest to the average). Consequently, $\mathcal{E}({\bf M})$ and $\mathcal{D}({\bf M},{\bf M}^{\prime})$ read
@@
while the formula for the cosine remains unaltered. This normalisation accounts for the specificity of the structural features of the protein, sets a characteristic scale for the value of scalar product, and enables a fair comparison between mapping pairs defined on molecules of different size.
Figure 2: Panels (a) to (c): high, random, and low norm mappings, respectively. Panels (d) and (e): parallel and orthogonal mappings, respectively. Figure reproduced from: Menichetti et al. 34, “A journey through mapping space: characterising the statistical and metric properties of reduced representations of macromolecules”, Eur. Phys. J. B 94, 204 (2021).
The norm of a mapping is a measure of its “compactness”: it was observed, in fact, that mappings with high value of the norm select groups of atoms that are very close to each other, hence corresponding to coarse pictures in which a relatively compact region of the molecule is represented with high resolution, while the remainder is largely discarded. In contrast, low-norm mappings correspond to very sparse selections, in which the retained atoms are maximally distant one from the other compatibly with their number and the properties of the molecule. Illustrative examples of these cases are reported in Fig. 2 (panels $a$ to $c$).
@@ 2.4 Summary of the EXCOGITO tools
At present, EXCOGITO contains the following subprograms:
  • measure: the user provides a mapping to EXCOGITO in the form of a text file (a prototype is available in the examples) and the associated mapping entropy $S_{map}^{\beta}$ (Eq. 14) is computed;
  • measure_kl: the Kullback-Leibler version of task measure;
  • random: generation of n_mappings (see Tab. 2) and measurement of the corresponding values of $S_{map}$. This task is useful when one wants to compare the values of mapping entropy of optimal mappings to those of coarse-grained representations randomly drawn from the mapping space;
  • random_kl: the Kullback-Leibler version of task random;
  • optimize: a mapping optimization run that produces $K$ local minima of the mapping entropy in the space of coarse-grained mappings. The number of minima $K$ has to be lower or equal to the number of CPU cores of the employed architecture, since each core performs a single minimization. The algorithm employed for the optimization is Monte Carlo simulated annealing: at each step, the current mapping $\bf M$} is slightly modified into a new one $\bf M$$^{\prime}$ by replacing an atom with another one that was not part of $\bf M$}. Such move is accepted or rejected with a probability given by a Metropolis criterion:
    $W({\bf M}\rightarrow{\bf M}^{\prime})=\text{min}\left[1,e^{(S_{map}({\bf M})-S_{map}({\bf M}^{\prime}))/T}\right].$ (24)
    where the simulated annealing temperature $T$ experiences an exponential decay in time dictated by:
    $T(i)=T_{0}\ e^{-i/\nu}.$ (25)
@@
  • optimize_kl: analogous to optimize, but using the Kullback-Leibler version of the mapping entropy ($S_{map}$, Eq. 12). More specifically, the user provides EXCOGITO with a set of atomistic configurations $\bf r$} (such as a MD trajectory), together with the associated, non-uniform probabilities $p({\bf r})$. A further clustering on this set of microstates partitions the conformational space in CG macrostates, each one having an associated probability given by the number of frames in the cluster. For each microstate, $p_{r}({\bf r})\ln\left(p_{r}({\bf r})/\bar{p}_{r}({\bf r})\right)$ measures the discrepancy between its probability and the smeared one;
  • norm: given a mapping and a trajectory, the time-evolution of the squared mapping norm (Eq. 22) is calculated. The value of the atomistic coordination number (Eq. 21) is chosen as the one calculated over the first structure provided in input;
  • cosine: given two mappings and a trajectory, the time-evolution of the cosine (Eq. 19) between them is calculated;
  • distance: given a set of n_mappings coarse-grained mappings and a given configuration of the molecule, the distance matrix between them is computed using Eq. 23. Such matrix can be employed for several purposes, such as the calculation of the sketch maps as in Ref. 34.
3 Previous applications of EXCOGITO
In this section we report a selection of the applications of EXCOGITO from previous works. First we address the calculation and minimisation of the mapping entropy; subsequently we discuss the usage of the mapping space metric tools.
@@ 3.1 Applications of the mapping entropy tools
Figure 3: Structure of Tamapin colored according to the probability of preserving an atom. Figure reproduced from: Giulini et al. 31, “An Information-Theory-Based Approach for Optimal Model Reduction of Biomolecules”, J. Chem. Theory Comput. 2020, 16, 11, 6795–6813 31.
The work by Giulini and coworkers  31 describes the first application of the mapping entropy optimisation workflow to three markedly different proteins, namely the tamapin toxin, adenylate kinase, and $\alpha$-11 antitrypsin. Upon simulating these molecules in explicit solvent, the distribution of values of $S_{map}^{\beta}$ is calculated on $500$ randomly selected mappings (random subcommand). Then, several optimisations are run (optimize subcommand), resulting in mappings that correspond to local minima of $S_{map}^{\beta}$. Upon averaging over these solutions, it is evident how the mapping entropy optimisation assigns an uneven level of detail to the structures, with some amino acids that are retained more often than others. In all the three considered cases, the retained amino acids are heavily involved in the biological role of the protein, and in particular in the binding to another molecule. This is a consequence of the fact that, in simulations performed in absence of the substrate, amino acids involved in the binding tend to correlate with important energetic fluctuations at the level of the whole protein, and feature high conformational variability. These two characteristics determine a higher chance for these residues to be retained in an optimal representation: in fact, the knowledge of their position and arrangement provides a better picture of the system as a whole at a lower resolution level, in comparison to other residues whose structural properties are less informative.
Fig. 3 shows an example of this behavior for the tamapin toxin: the minimisation of the mapping entropy for various degrees of resolution (i.e. for different numbers of retained atoms) consistently leads to the conservation of terminal atoms of ARG6 and ARG13, which are the two amino acids responsible for the binding to the toxin’s substrate, the SK2 calcium-activated potassium channel.

### v2

@@ 2.2 Mapping entropy and related EXCOGITO tools
EXCOGITO enables the identification of the maximally informative CG representation of a system via the optimize and optimize_kl tools, which respectively build on the equilibrium and KL mapping entropy estimators reported in Eqs. 14 and  15. Both protocols minimize $S_{map}$ by relying on a Monte Carlo simulated annealing (SA) approach, gradually pushing a stochastic exploration of the mapping space of the system to visit CG representations characterized by a low information loss, see Sec. 2.4 for all technical details. More specifically, starting from an initial selection of $N$ sites of the macromolecule, optimize and optimize_kl perform a sequence of Monte Carlo moves that, at each step, propose a swap between a retained and neglected atom in the CG mapping, hence working at a fixed degree of coarse-graining. The moves are accepted according to a Metropolis-like criterion that employs $S_{map}$ as a cost function, and in which the associated “temperature” is exponentially decreased in the course of the simulation, driving the sampling, after an initial transient, to converge towards a local minimum of the mapping entropy. A single SA run performed with one of the two methods thus enables the identification of one of the sought-for maximally informative CG representations that can be employed to describe the system of interest.
The next necessary step in this analysis is to account for the fact that the manifold of solutions to the optimization problem can have a rather complex structure. Indeed, given the intricacy of the network of interactions among the system’s microscopic constituents, it is reasonable to expect a rugged landscape of information loss throughout the mapping space, exhibiting a whole ensemble of more or less degenerate local minima either living in relatively flat basins or being widely separated by high $S_{map}$ barriers. None of these minima can a priori be preferred over another; rather, in order to gather a full picture of the link between resolution reduction and information content, one needs to simultaneously consider a pool of solutions minimising the mapping entropy. As we will discuss in the following, it is precisely the pattern that emerges from the analysis of the whole ensemble of such optimized solutions that enables the extraction of nontrivial information about the system and its function.
With reference to the software, the optimize and optimize_kl tools of EXCOGITO enable the user to tackle the simultaneous extraction of several (local) minima of the mapping entropy through multiple SA runs that are performed in parallel, see Sec. 2.4 for all technical details.
@@ 2.3 Metric in the coarse-grained mapping space and related EXCOGITO tools
Equations 14 and 15 allow us to calculate the mapping entropy associated to a given mapping, and their optimisation leads to the identification of mappings that entail the largest amount of information about the system for a given number of retained atoms. It is also instructive, however, to broaden the perspective on mappings themselves, and to investigate their properties from a purely structural perspective.
More specifically, given a mapping as the selection of a particular subgroup of elements (the retained atoms) from a set (the whole molecule), we can ask ourselves questions about the total number of mappings, the amount of them sharing the same qualitative and quantitative features, and the relationship that exists between mapping groups with given features and the structural properties of the system on which the selection takes place. These questions have been addressed in various recent works 35, 34, 36; here, we only report those results obtained by some of us 34 through the application of the methods implemented in EXCOGITO.
In order to assess even the simplest properties of the mapping space associated to a given protein, one needs to define basic quantitative instruments, for example to determine how different a given mapping is from another. To this end, some of us introduced a notion of norm, cosine, and distance between CG mappings, which only make use of the structural features of the biomolecule of interest. More specifically, we define the scalar product $\langle{\bf M},{{\bf M}^{\prime}}\rangle$ between two mappings $\bf M$} and $\bf M$$^{\prime}$ as:
@@
where $\chi_{{\bf M},j}$ is the mapping function as defined in Eq.7, $r_{ij}$ is the distance between atoms $i$ and $j$, and $\sigma$ is a parameter that tunes the amplitude of the Gaussian employed to calculate the coupling $J_{ij}=e^{-r^{2}_{ij}/4\sigma^{2}}$ between them. In the work in which it was introduced, this parameter was set to $\sigma=0.19$ nm, however its value can be tuned according to the specific application of interest.
From Eq. 17 we can calculate the (squared) norm of a mapping as
@@
In Ref. 34 the norm and the distance were rescaled by a function of the atomistic coordination number:
@@
calculated over a specific molecular configuration (e.g. the initial frame of an MD trajectory, or the frame closest to the average conformation of the system). Consequently, $\mathcal{E}({\bf M})$ and $\mathcal{D}({\bf M},{\bf M}^{\prime})$ read
@@
while the formula for the cosine remains unaltered. This normalisation accounts for the specificity of the structural features of the protein, sets a characteristic scale for the value of scalar product, and enables a fair comparison between mapping pairs defined on molecules of different size 34.
Figure 2: Panels (a) to (c): examples of low, intermediate, and high norm mappings, respectively, for adenylate kinase. Panels (d) and (e): parallel and orthogonal mappings, respectively. Figure reproduced from: Menichetti et al. 34, “A journey through mapping space: characterising the statistical and metric properties of reduced representations of macromolecules”, Eur. Phys. J. B 94, 204 (2021). This image is licensed under $C$ BY 4.0 http://creativecommons.org/licenses/by/4.0.
The norm of a mapping, or equivalently its square in Eq. 22, is a measure of the mapping’s “compactness”: it was observed, in fact, that CG representations with high value of the norm select groups of atoms that are very close to each other, hence corresponding to coarse pictures in which a relatively compact region of the molecule is represented with high resolution, while the remainder is largely discarded. In contrast, low-norm mappings correspond to very sparse selections, in which the retained atoms are maximally distant one from the other compatibly with their number and the properties of the molecule. Illustrative examples of these cases are reported in Fig. 2 (panels $a$ to $c$).
@@ 2.4 Summary of the EXCOGITO tools
At present, EXCOGITO contains the subprograms listed and described hereafter. The organisation of these methods and their relationship, input, and output is graphically illustrated in Fig. 3. Technical aspects of the execution of EXCOGITO are provided in Sect. 6.1 of the Appendix, where we list the command line syntax of the various methods in Tab. 2 as well as the parameters required in input in Tab. 3.
  • measure: the user provides in input to EXCOGITO a series of configurations extracted from an equilibrium MD simulation of the system of interest, the associated energies, and a specific choice of the CG mapping in the form of a text file—a prototype is available in the examples. The associated mapping entropy $\tilde{S}_{map}$ is then computed according to Eq. 14;
  • measure_kl: Kullback-Leibler-based routine for the computation of the mapping entropy via Eq. 15. Here, the user provides EXCOGITO with a set of atomistic configurations, the associated probabilities, as well as a CG mapping;
  • random: generation of n_mappings random CG mapping of the system with a fixed number of cgnum sites (see Tab. 3) and measurement of the corresponding values of $\tilde{S}_{map}$. This task is useful, e.g., when one wants to compare the values of $\tilde{S}_{map}$ of optimal mappings to those of CG representations that are randomly drawn from the mapping space;
  • random_kl: Kullback-Leibler version of task random, making use of the $\hat{S}_{map}$ estimator;
  • optimize: optimization task that produces $K$ local minima of the mapping entropy $\tilde{S}_{map}$ in the space of the possible CG representations of the system that retain cgnum sites (see Tab. 3). The number of minima $K$ has to be lower or equal to the number of CPU cores of the employed architecture since each core performs a single minimization. The algorithm employed for the optimization is Monte Carlo simulated annealing; at each step, the current mapping $\bf M$} is slightly modified into a new one $\bf M$$^{\prime}$ by replacing a retained atom with another one that was not part of $\bf M$}. Such a move is accepted or rejected with a probability given by a Metropolis criterion:
    $W({\bf M}\rightarrow{\bf M}^{\prime})=\text{min}\left[1,e^{(\tilde{S}_{map}({\bf M})-\tilde{S}_{map}({\bf M}^{\prime}))/T}\right],$ (24)
    in which the simulated annealing temperature $T$ experiences an exponential decay in time dictated by
    $T(i)=T_{0}\ e^{-i/\nu},$ (25)
@@
  • optimize_kl: analogous to optimize, but using the Kullback-Leibler version of the mapping entropy $\hat{S}_{map}$;
  • norm: given a CG representation and a trajectory, the time-evolution of the squared mapping norm in Eq. 22 is calculated. The value of the atomistic coordination number (Eq. 21) is chosen as the one calculated over the first structure provided in input;
  • cosine: given two CG representations and a trajectory, the time-evolution of the cosine (Eq. 19) between them is calculated;
  • distance: given a set of n_mappings (see Tab. 3) CG representations and a given configuration of the molecule, the distance matrix between them is computed using Eq. 23. Such matrix can be employed for several purposes, such as the calculation of the sketch maps as in Ref. 34.
Figure 3: Scheme of the EXCOGITO methods, inputs, and outputs. The starting point is a sample of molecular configurations, or a given structure thereof; these can be obtained from an equilibrium MD trajectory, but other sources (e.g. non-equilibrium MD simulations, ensembles of experimental structures) are possible. Adding the potential energy (in the case of equilibrium sampling) or the probability associated with each frame one can estimate the value of the mapping entropy of a given CG representation, generate a number of random mappings, and eventually optimise $S_{map}$ through simulated annealing. Alternatively, the trajectory can be analysed through the computation of the norm of a given CG mapping and/or the scalar product between two or more mappings throughout the frames.
3 Previous applications of EXCOGITO and software performance
In this section, we report a selection of the applications of EXCOGITO from previous works. We begin by addressing the calculation and minimisation of the mapping entropy; subsequently, we discuss the usage of the mapping space metric tools; lastly, we provide information on the performance of the software in a few selected cases, as indicative guidelines on the running time needed for performing the tasks of EXCOGITO.
@@ 3.1 Applications of the mapping entropy tools
Figure 4: Structure of Tamapin colored according to the probability of preserving an atom in the pool of EXCOGITO optimised mappings. Figure reproduced from: Giulini et al. 31, “An Information-Theory-Based Approach for Optimal Model Reduction of Biomolecules”, J. Chem. Theory Comput. 2020, 16, 11, 6795–6813, https://pubs.acs.org/doi/full/10.1021/acs.jctc.0c00676. Further permissions related to this figure should be directed to the ACS.
The work by Giulini and coworkers  31 describes the first application of the mapping entropy optimisation workflow to three markedly different proteins, namely the tamapin toxin, adenylate kinase, and $\alpha$-11 antitrypsin. Upon simulating these molecules in explicit solvent, the distribution of values of $\tilde{S}_{map}$ is calculated on $500$ randomly selected mappings (random subcommand). Then, several optimisations are run (optimize subcommand), resulting in mappings that correspond to local minima of $\tilde{S}_{map}$. Upon averaging over these solutions, it is evident how the mapping entropy optimisation assigns an uneven level of detail to the structures, with some amino acids that are retained more often than others. In all the three considered cases, the retained amino acids are heavily involved in the biological role of the protein, and in particular in the binding to another molecule 31. This is a consequence of the fact that, in simulations performed in absence of the substrate, amino acids involved in the binding tend to correlate with important energetic fluctuations at the level of the whole protein, and feature high conformational variability. These two characteristics determine a higher chance for these residues to be retained in an optimal representation: in fact, the knowledge of their position and arrangement provides a better picture of the system as a whole at a lower resolution level, in comparison to other residues whose structural properties are less informative.
Fig. 4 shows an example of this behavior for the tamapin toxin: the minimisation of the mapping entropy for various degrees of resolution (i.e. for different numbers of retained atoms) consistently leads to the conservation of terminal atoms of ARG6 and ARG13, which are the two amino acids responsible for the binding to the toxin’s substrate, the SK2 calcium-activated potassium channel 31.
@@ 3.2 Applications of the mapping space metric tools
3.3 Performance and optimisation
The runtime of EXCOGITO largely depends on several factors. First, the various tasks have very different computational requirements, as some of them consist of rapid one-time calculations (e.g., individual computations of norm, cosine, or measure, such tasks being listed in order of increasing effort), while others (most notably optimize and optimize_kl) are iterative, and thus intrinsically much more time-consuming. Furthermore, also the system of interest affects the performance, most notably through its size (number of atoms/residues of the protein) and the number of configurations passed as input to the software. Finally, additional ingredients that influence the performance of EXCOGITO are, quite naturally, the architecture and efficiency of the machine it is run on, as well as the compilation options. The latter, in particular, can considerably speed up the calculations; to this end, a few general guidelines to achieve this boost through compiler optimisation are provided in the README of the software, thereby leaving the users the opportunity to improve the efficiency in the most appropriate manner for their equipment.
From the preceding discussion, it follows that a comprehensive assessment of the computational effort required to perform all the possible EXCOGITO tasks, further covering different combinations of the input parameters, would be difficult to achieve. To help the users roughly estimate the computational and time resources needed for performing their analyses, we here focus our attention on the computing time associated with the fundamental building block of EXCOGITO, namely the measure tool calculating the $\tilde{S}_{map}$ associated the choice of a specific CG representation of the system, see Eq. 14. Critically, this task lies at an intermediate level of computational effort among all the EXCOGITO ones.
The average runtime necessary for performing a single estimate of $\tilde{S}_{map}$ for a random CG mapping via measure is presented in Table 1 for three biomolecules of increasing size, namely tamapin, adenylate kinase, and $\alpha$-11-antitrypsin.31 In each of the three cases, we fix the number of CG sites to be equal to the number of residues in the protein and resort to the standard clustering criterion implemented in EXCOGITO for extracting the CG macrostates (criterion = 0 in Sec. 6.2), reporting results for different numbers of configurations provided in input to the software. For the largest number of frames ($10^{4}$), the computational times required to perform the calculations via the standard clustering method on the three systems are further compared to those obtained by relying on the fast clustering strategy (criterion = 3, stride = 3); as discussed in Sec. 6.2, the latter can be employed only provided that a time-continuous MD trajectory of the system is available. All computations were performed on a single core of an Intel Xeon-Gold 5118 processor, compiling the software in a plain vanilla setup without any optimisation. Overall, in the case of the standard clustering, we observe that for a fixed number of input configurations the runtime of EXCOGITO scales linearly with the number of (heavy) atoms in the system. On the other hand, for a fixed number of constituent atoms, the computational time scales quadratically with the number of frames. Finally, we note that employing the fast clustering method can speed up $S_{map}$ roughly by a factor of 10. As anticipated, we stress that the variables that can affect the program performance are many, hence the real runtime can vary with respect to the estimates reported in Table 1.
Protein $n_{at}$ $N_{CG}$ CPU time $(s)$
$f=1000$ $f=2500$ $f=5000$ $f=10000$ F.C.
TAM 230 31 1 5.8(4) 25(2) 105(7) 12(1)
AKE 1656 214 7.1(7) 41(3) 160(4) 631(8) 74(1)
AAT 2956 372 12.7(6) 72(2) 279(2) 1102(4) 130(2)
Table 1: Average CPU time (in seconds) needed to run a single $S_{map}$ estimate via the EXCOGITO measure tool for three proteins with an increasing number of constituent atoms $n_{at}$—namely tamapin (TAM), adenylate kinase (AKE), and $\alpha$-11-antitrypsin (AAT).31 The number of retained sites $N_{CG}$ was in each case set equal to the number of residues in the molecule, and the computational time reported was calculated as an average over 2020 randomly extracted CG mappings of the system. We display the results associated with the standard clustering method of EXCOGITO (see Sec. 6.2) for different numbers of input frames $f$ employed in the calculations, as well as, for $f=10^{4}$, the runtime one achieves by relying on the fast clustering strategy (column “F.C.”, see Sec. 6.2). Computations were performed on a single core of an Intel Xeon-Gold 5118 processor.